OpenStudy (anonymous):

Solve each system by elimination: 3x+2y=-18 5x+6y=-30

OpenStudy (anonymous):

ok- remember how in the other one we had to find where we could get a negative and positive value for one of the variables in order to eliminate it? try looking at the x and y values in each equation to see if you can multiply one to get the positive and negative values.

OpenStudy (anonymous):

umm i think 12y and 6y? im not sure,correct me if im wrong

OpenStudy (anonymous):

???

OpenStudy (anonymous):

the 2y and 6y- is that what you meant? (sorry i had to leave my computer) the 2y can be multiplied by -3 to get -6y, which, when added to 6y, will eliminate it.

OpenStudy (anonymous):

OH GOSH! sorry I messed up the problem its suppose to be 3x+12y=-18 5x+6y=-30

OpenStudy (anonymous):

ok. haha. yes- then it would be the 12y and the 6y. great job! do you know what you need to do for the next step?

OpenStudy (anonymous):

honestly no.arent you suppose to divide them?...i think

OpenStudy (anonymous):

that is how you would find the multiplier (the 3 from the last problem we did)

OpenStudy (anonymous):

so then would we have to divide?

OpenStudy (anonymous):

12/6 would give us 2, which is what we need to multiply the second equation by to get 12y in that equation. im doing a really poor job of explaining this, i'm sorry.

OpenStudy (anonymous):

nah dont worry its ok.im just really thankful that your actually helping me.it helps me better when you draw it though.pictures help me alot better

OpenStudy (anonymous):

ok. pictures it is.

OpenStudy (texaschic101):

3x + 12y = - 18 5x + 6y = - 30 --->(-2)5x + 6y = - 30 --------------- 3x + 12y = - 18 -10x - 12y = 60 --------------- -7x = 42 -7x/-7 = 42/-7 x = - 6 now sub -6 in for y in either of the equations. 3x + 12y = - 18 3(-6) + 12y = - 18 - 18 + 12y = - 18 12y = - 18 + 18 y = 0/12 = 0 check... 5x + 6y = - 30 5(-6) + 6(0) = - 30 - 30 + 0 = - 30 - 30 = - 30 (correct) ANSWER : x = -6 and y = 0

OpenStudy (anonymous):

 notice how in this one, if you were to add the equations together, neither variable would be eliminated. thus, we use subtraction to eliminate the variable: